A local civic theater has 22 seats in the first row and 21 rows in all. Each successive row contains 3 additional seats. How many seats are in the civic theater?
a.) 1092
b.) 1010
c.) 1070
d.) 790

Answer by Boreal(15235)

sum of the 21 rows, s21=21/2(22+sn), where the last row, 20th back, has 60 more seats or 82.
Therefore, s21=21/2(22+82)=21/2(104)=1092 seats. The formula takes the first and the last and uses that as an average, multiplying by the number of rows.
22+25+28+31=116
34+37+40+43=154
46+49+52+55=202
58+61+64+67=250
70+73+76+79=298
82==========82
That equals 1092.